The graph of the system of equations −5x+y=1 and 20x+4y=4 consists of what?
1.two distinct parallel lines
2.one line
3.two lines intersecting where x=1
4.two lines intersecting where y=1
two lines intersecting where y=4

Answer by Cromlix(4381)

Sort both equations into y = mx + c
-5x + y = 1
y = 5x + 1
20x + 4y + 4
4y = -20x + 4
Divide both sides by 4
y = -5x + 1
Solve:
y - 5x = 1 .....1
y + 5x = 1......2
Add 1 + 2
2y = 2
y = 1
Substitute y = 1 in Eq(1)
1 - 5x = 1
-5x = 0
x = 0
Two lines intersecting at y = 1 (4)
Hope this helps.
:-)